4. Write the augmented matrix of the system shown below: (5x – y - z = 0 x + y = 5 2x – 3z = 2
Answer #1

Given equations: 5x-y-z=0 x+y=5 2x-3z=2 The augumented matrix is: 5 -1 1 1 0 0 5 -3 2 2 0